OpenStudy (anonymous):

can you help with this derivative? d/dx∓(sqrt(5/2-3/2*x²)) what's the right differentiation rule to use?

OpenStudy (anonymous):

\[\frac{ d }{ dx }∓\sqrt{\frac{ 5 }{ 2 }-\frac{ 3x² }{ 2 }}\]

sam (.sam.):

What's with \(∓\)? is it =?

OpenStudy (anonymous):

it's plusminus

OpenStudy (anonymous):

I think I got it. I'll close this now

sam (.sam.):

mkay

OpenStudy (anonymous):

mmkay?